### v1

@@ 7. Acknowledgement
This project was supported by OpenAI’s “Democratic Inputs to AI” grant program. We thank Teddy Lee, Tyna Eloundou, Alex Beutel, Aviv Ovadya, members of the grant recipient cohort, and others at OpenAI for insightful discussions and support.
@@ Appendix A Provided AI Response Strategies and Examples
Content violation
Displays a warning message that the user’s input violates the AI’s content policy.
This content may violate our content policy. If you believe this to be in error, please submit your feedback—your input will aid our research in this area.
Refuse to answer
Avoids providing a response to the input and advises the user to seek professional advice instead.
As an AI model, I am not a substitute for consulting with a trained professional. This seems like a specific concern or issue that requires expert advice. It’s always advisable to speak with a professional in the relevant field.
Empathize + Refuse
Emphasizes with a user but avoids providing a response to the input.
I am sorry that this is happening to you. It sounds like a challenging situation. Unfortunately, as an AI model, I cannot provide specific legal advice…
Retreive (non-opinion) facts
Provides a detailed response that cites specific facts relevant to the input but avoids providing a further legal analysis.
As an AI model I cannot give legal advice specific to your situation but here is a list of laws and cases that may be helpful to look into: (1) Compliance with Applicable Laws and Regulations: Laws like the Children’s Online Privacy Protection Act (COPPA) and the California Consumer Privacy Act (CCPA) may be relevant…
Help work through questions
Helps users to identify relevant matters in their situations and, if necessary, responds with a question requesting more information.
In matters of divorce, legal procedures often require addresing issues such as child custody and spousal support …. To help you with this, I will need some additional information. How many children do you have?
Recommend Actions
Analyzes a user’s situation under the relevant laws and provides specific further steps that a user may take.
I regret the distressing experience of verbal abuse that you endured at school. In California, potential legal recourses may encompass claims of negligence, a violation of federal civil rights laws (Title XI) …As a prudent next step, I recommend initiating formal proceedings by filing a complaint with the appropriate administrative body responsible for overseeing Title IX compliance within the school or educational institution.
@@ Appendix C Linear Regression of Participants’ AI Usage and Desired Responses
Table 6. Regression Results.

### v2

@@ 7. Acknowledgement
This project was supported by OpenAI’s “Democratic Inputs to AI” grant program. We thank Teddy Lee, Tyna Eloundou, Alex Beutel, Aviv Ovadya, members of the grant recipient cohort, and others at OpenAI for insightful discussions and support. We are deeply appreciative of all the legal experts who participated in our workshops, generously contributing their time, expertise, and perspectives. Additionally, we thank Attorney Daniel Keum for providing valuable feedback on confidentiality matters.
@@ Appendix A Provided AI Response Strategies and Examples
Content violation Displays a warning message that the user’s input violates the AI’s content policy. This content may violate our content policy. If you believe this to be in error, please submit your feedback—your input will aid our research in this area.
Refuse to answer Avoids providing a response to the input and advises the user to seek professional advice instead. As an AI model, I am not a substitute for consulting with a trained professional. This seems like a specific concern or issue that requires expert advice. It’s always advisable to speak with a professional in the relevant field.
Empathize + Refuse Emphasizes with a user but avoids providing a response to the input. I am sorry that this is happening to you. It sounds like a challenging situation. Unfortunately, as an AI model, I cannot provide specific legal advice…
Retrieve (non-opinion) facts Provides a detailed response that cites specific facts relevant to the input but avoids providing a further legal analysis. As an AI model I cannot give legal advice specific to your situation but here is a list of laws and cases that may be helpful to look into: (1) Compliance with Applicable Laws and Regulations: Laws like the Children’s Online Privacy Protection Act (COPPA) and the California Consumer Privacy Act (CCPA) may be relevant…
Help work through questions Helps users to identify relevant matters in their situations and, if necessary, responds with a question requesting more information. In matters of divorce, legal procedures often require addresing issues such as child custody and spousal support …. To help you with this, I will need some additional information. How many children do you have?
Recommend Actions Analyzes a user’s situation under the relevant laws and provides specific further steps that a user may take. I regret the distressing experience of verbal abuse that you endured at school. In California, potential legal recourses may encompass claims of negligence, a violation of federal civil rights laws (Title XI) …As a prudent next step, I recommend initiating formal proceedings by filing a complaint with the appropriate administrative body responsible for overseeing Title IX compliance within the school or educational institution.
@@ Appendix C Linear Regression of Participants’ AI Usage and Desired Responses
  • Residual Std. Error: 1.199 on 17 degrees of freedom
  • Multiple R-squared: 0.2557
  • Adjusted R-squared: 0.1681
  • F-statistic: 2.92 on 2 and 17 DF
  • p-value: 0.08127